# Efficient absorbing boundary conditions for conservation laws: Application to the piston problem of gas dynamics

### v1

@@ 3.1 A relation between the RM and the far-field damping operator
$=\mathbf{q}^{*}+\Delta t\frac{1-\Gamma(x)}{\Delta t}(\overline{\mathbf{q}}-{\mathbf{q}})$ (12b)
@@
$=\mathbf{R}\begin{pmatrix}\frac{\tilde{\Gamma}(x;\Delta t)\lambda_{1}}{\lambda_{1}}&0&\cdots&0\\ 0&\frac{\tilde{\Gamma}(x;\Delta t)\lambda_{2}}{\lambda_{2}}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&\frac{\tilde{\Gamma}(x;\Delta t)\lambda_{N}}{\lambda_{N}}\end{pmatrix}\mathbf{R}^{-1}(\overline{\mathbf{q}}-\mathbf{q})$ (15b)
$=\mathbf{R}{\bf\Lambda_{D}}(x;\Delta t)\mathbf{R}^{-1}(\overline{\mathbf{q}}-\mathbf{q})=\mathbf{D}(x;\Delta t)(\overline{\mathbf{q}}-\mathbf{q}),$ (15c)
@@
$d_{i}^{\text{RM}}(x;\Delta t)=\frac{\tilde{\Gamma}(x;\Delta t)}{\lambda_{i}}=\frac{1-\Gamma(x)}{\Delta t\lambda_{i}}.$ (16)
Moreover, the strength of left-going waves ($\lambda_{j}<0$) will be amplified according to Equation (6), since $d_{j}^{RM}\leq 0$. The inversely proportional relation between $d_{i}$ and $\Delta t$ is consistent with the increase in damping severity as the time step size decreases using the RM observed in the literature (see, e.g., [36]).
@@ 3.2 The RM with matrix-valued weight function (RM-M)
We propose a simple extension of the RM such that it resembles the directional behavior of Karni’s damping far-field operator. We achieve this by allowing the weight function ${\mathbf{\Gamma}}$ to be matrix-valued and defined in terms of $\mathbf{R}$ (and therefore $\mathbf{q}$ in the nonlinear case). The process is described in Algorithm 2 and justified as follows. Proceeding analogously to the previous section, the relaxation step in Algorithm 2
$\mathbf{q}^{[n+1]}={\bf\Gamma}(x)\mathbf{q}^{*}+(\mathbf{I}-{\bf\Gamma}(x))\overline{\mathbf{q}},$ (17)
@@
$\partial_{t}\mathbf{q}=\tilde{\bf{\Gamma}}(x,\mathbf{q};\Delta t)(\overline{\mathbf{q}}-\mathbf{q}),$ (18)
@@
$\tilde{\bf{\Gamma}}(x,\mathbf{q};\Delta t)=\frac{1}{\Delta t}(\mathbf{I}-{\bf\Gamma}(x))=\mathbf{R}\begin{pmatrix}\frac{1-\Gamma_{1}(x)}{\Delta t}&0&\cdots&0\\ 0&\frac{1-\Gamma_{2}(x)}{\Delta t}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&\frac{1-\Gamma_{N}(x)}{\Delta t}\end{pmatrix}\mathbf{R}^{-1}.$ (19)
@@
$d_{i}^{\text{RM-M}}(x)=\frac{1-\Gamma_{i}(x)}{\Delta t\lambda_{i}},$ (20)
@@
If $\mathbf{R}$ is independent of $\mathbf{q}$, the matrix-valued weight function ${\mathbf{\Gamma}}$ does not have to be computed at each time step in Algorithm 2. Additionally, if the far-field state $\overline{\mathbf{q}}$ is constant, we can define the translated function $\mathbf{v}=\mathbf{q}-\overline{\mathbf{q}}$ that satisfies the conservation law $\partial_{t}\mathbf{v}+\partial_{x}\mathbf{f}(\mathbf{v}+\overline{\mathbf{q}})=0$. This system may be solved with the far-field state $\bf 0$ and $\mathbf{q}$ recovered afterward. In this case, we have
$||\mathbf{v}^{[n+1]}||_{2}\leq||{\mathbf{\Gamma}}\mathbf{v}^{*}||_{2}\leq||{\mathbf{\Gamma}}||_{2}||\mathbf{v}^{*}||_{2}\leq||\mathbf{v}^{*}||_{2},$ (21)
@@
$\mathbf{q}^{*}\leftarrow$ Evolve a discretization of (1) over a time step $\Delta t$ with initial data $\mathbf{q}^{[n]}$triangleright Hyperbolic step
@@
${\bf\Lambda}_{\mathbf{\Gamma}}(x)\leftarrow$diagonal matrix-valued function with entries $\Gamma_{i}(x)$
${\bf{\Gamma}}(x)\leftarrow\mathbf{R}{\bf\Lambda}_{\mathbf{\Gamma}}(x)\mathbf{R}^{-1}$ where $\mathbf{R}$ contains the right eigenvectors of $\mathbf{f}^{\prime}(\mathbf{q})$
$\mathbf{q}^{[n+1]}\leftarrow{\bf\Gamma}(x)\mathbf{q}^{*}+(\mathbf{I}-{\bf\Gamma}(x))\overline{\mathbf{q}}$triangleright Relaxation step
@@ 3.3 On attaining higher order of accuracy in time with the RM
Let us consider Algorithm 2 for the linear problem (2), where the matrix-valued weight function $\mathbf{\Gamma}$ is independent of $\mathbf{q}$. Then, the relaxation step (17) could be written as a high-order discretization of a system $\partial_{t}\mathbf{q}=\mathcal{R}(g(\mathbf{\Gamma}(x),\Delta t))\,(\overline{\mathbf{q}}-\mathbf{q})$, where $\mathcal{R}$ is the stability function of some Runge-Kutta scheme [10]. Moreover, we can interpret it as the exact solution to the problem
@@
${\bf\Gamma}(x;\Delta t)=\exp({-\Delta t\mathbf{W}(x;\Delta t)}).$ (23)
Since ${\bf\Gamma}=\mathbf{R}{\bf\Lambda}_{\mathbf{\Gamma}}\mathbf{R}^{-1}$, then $\mathbf{W}=\mathbf{R}{\bf\Lambda_{W}}\mathbf{R}^{-1}$, where
@@
with $\Gamma_{i}$ and $w_{i}$ denoting the main diagonal entries of the matrices ${\bf\Lambda}_{\mathbf{\Gamma}}$ and ${\bf\Lambda_{W}}$ respectively. Thus, the $i$-th wave is damped at the rate
$d_{i}^{\mathbf{W}}(x)=\frac{w_{i}(x)}{\lambda_{i}}=\frac{-\ln\left(\Gamma_{i}(x)\right)}{\Delta t\lambda_{i}},$ (25)
and will be completely absorbed as $\Gamma_{i}\to 0$, i.e., at the boundary of the sponge layer $x_{\infty}$. This further justifies the approaches proposed by Seng [36] and Wasistho et al. [39] where, through different arguments, $\Gamma$ is substituted by $\Gamma^{C\Delta t}$ to reduce the damping sensitivity to the time step size.
@@
and a natural approach to get a higher-order accurate in time approximate solution to (26) would be to use the RM in the context of high-order operator splitting. For instance, we could use Strang splitting as described in Algorithm 3. However, in this setting, both splitting techniques are equivalent up to the choice of the scalar weight functions $\Gamma_{i}$. This can be seen as follows: Assuming that the solution is initialized as $\overline{\mathbf{q}}$ inside the sponge layer, the first relaxation step at $t^{[0]}=0$ does not alter the solution. Similarly, the second relaxation step in Algorithm 3 at the final time does not affect the solution in the computational domain, since it is local and no hyperbolic step is taken afterward. Finally, the following relation holds between $\mathbf{q}^{**}$ (computed at time $t^{[n]}$) and $\mathbf{q}^{*}$ (computed at time $t^{[n+1]}$)
$\mathbf{q}^{*}=\mathbf{\Gamma}\mathbf{q}^{[n+1]}+(\mathbf{I}-\mathbf{\Gamma})\overline{\mathbf{q}}=\mathbf{\Gamma}(\mathbf{\Gamma}\mathbf{q}^{**}+(\mathbf{I}-\mathbf{\Gamma})\overline{\mathbf{q}})+(\mathbf{I}-\mathbf{\Gamma})\overline{\mathbf{q}}=\mathbf{\Gamma}^{2}\mathbf{q}^{**}+(\mathbf{I}-\mathbf{\Gamma}^{2})\overline{\mathbf{q}}.$ (27)
@@
On the other hand, since $\mathbf{\Gamma}$ depends on $\mathbf{q}$ in the nonlinear case, the last observation is no longer valid, and Algorithms 2 and 3 are not equivalent a priori. The effect of applying the RM as in Algorithm 3 in the nonlinear setting is evaluated numerically in Section 5.3.
@@
Compute the weight function ${\bf\Gamma}(x)$ as in Algorithm 2
@@
$\mathbf{q}^{**}\leftarrow$ Evolve a discretization of (1) over a time step $\Delta t$ with initial data $\mathbf{q}^{*}$triangleright Hyperbolic step
@@ 3.4 Nonlinear damping far-field operator (NDO)
We propose an alternative extension of the damping operator from Section 2.1 to hyperbolic systems of the form (1) with a general nonlinear flux function. Here, the change of variables $\mathbf{w}=\mathbf{R}^{-1}\mathbf{q}$ no longer allows us to decompose a nonlinear system of the form (1) into a system of scalar evolution equations. Instead, we consider the nonlinear characteristic variables $\boldsymbol{\alpha}=\mathbf{R}^{-1}\partial_{x}\mathbf{q}$ which, after some manipulations to System (1) (see [27, Ch.1]), yield
@@
$\partial_{t}\alpha_{i}+\partial_{x}\left(\lambda_{i}(\mathbf{q})\alpha_{i}\right)=\mathcal{I}(\mathbf{q},\boldsymbol{\alpha}),\quad 1\leq i\leq N,$ (28)
@@
where $\mathcal{I}(\mathbf{q},\boldsymbol{\alpha})$ denotes the nonlinear interactions between the $N$ characteristic fields. Although this is not a decoupled system of equations, we can add a damping term at the right-hand side, resembling the modification done in (5) for the linear equations (2), to obtain
@@
$\partial_{t}\alpha_{i}+\partial_{x}\left(\lambda_{i}(\mathbf{q})\alpha_{i}\right)=\mathcal{I}(\mathbf{q},\boldsymbol{\alpha})-d_{i}(x)\lambda_{i}(\mathbf{q})\alpha_{i},\quad 1\leq i\leq N.$ (29)
@@
After multiplying this system from the left by $\mathbf{R}$ and integrating from $x$ to +$\infty$+ ∞ (inverting the process carried out to obtain (28)), we get
$\partial_{t}\mathbf{q}(t,x)+\partial_{x}\mathbf{f}(\mathbf{q}(t,x))$ $=\int_{x}^{x_{\infty}}\mathbf{R}\boldsymbol{\Lambda}_{\mathbf{D}}(z){\mathbf{R}}^{-1}\partial_{z}\mathbf{q}(t,z)\,dz$ (30a)
$=\int_{x}^{x_{\infty}}\mathbf{D}(z,\mathbf{q})\partial_{z}\mathbf{q}(t,z)\,dz,$ (30b)
@@
where the artificial boundary at the end of the sponge layer is enforced through the assumption $\mathbf{q}(t,x)=\overline{\mathbf{q}}$ for $x\geq x_{\infty}$. If the damping rate functions $d_{i}$ are defined as in Section 2.1, the term (30b) vanishes outside of the sponge layer. Moreover, if the flux is linear and each $d_{i}$ is constant, the damping far-field operator of Section 2.1 is recovered.
@@
Some challenges that arise when dealing with the nonlinear damping term (30b) are its non-local nature and the presence of a nonconservative product. This term is significantly simplified if one can write $\mathbf{D}(z,\mathbf{q})=\partial_{\mathbf{q}}\mathbf{h}(z,\mathbf{q})$ for some function $\mathbf{h}\in C^{1}(\times^{N},^{N})$. Indeed, in that case, we would have
$\int_{x}^{x_{\infty}}\mathbf{D}(z,\mathbf{q})\partial_{z}\mathbf{q}(t,z)\,dz$ $=\int_{x}^{x_{\infty}}\partial_{\mathbf{q}}\mathbf{h}(z,\mathbf{q})\partial_{z}\mathbf{q}(t,z)\,dz=\int_{x}^{x_{\infty}}\frac{d}{dz}\mathbf{h}(z,\mathbf{q})-\partial_{z}\mathbf{h}(z,\mathbf{q})\,dz$ (31a)
$=\mathbf{h}(x_{\infty},\overline{\mathbf{q}})-\mathbf{h}(x,{\mathbf{q}(t,x)})-\int_{x}^{x_{\infty}}\partial_{z}\mathbf{h}(z,\mathbf{q}(t,z))\,dz.$ (31b)
@@
Nevertheless, the integrability conditions to guarantee the existence of such a function $\mathbf{h}$ do not hold for a general system and choice of $d_{i}$’s [32]. Thus, for the application studied in Section 5 we work with a discretization of the term (30b), described in Appendix A.2.

### v2

@@ 3.1 A relation between the RM and the far-field damping operator
$=\mathbf{q}^{*}+\Delta t\frac{1-\Gamma(x)}{\Delta t}(\overline{\mathbf{q}}-{\mathbf{q}}^{*})$ (18b)
@@
$=\mathbf{R}\begin{pmatrix}\frac{\tilde{\Gamma}(x;\Delta t)|\lambda_{1}|}{|\lambda_{1}|}&0&\cdots&0\\ 0&\frac{\tilde{\Gamma}(x;\Delta t)|\lambda_{2}|}{|\lambda_{2}|}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&\frac{\tilde{\Gamma}(x;\Delta t)|\lambda_{N}|}{|\lambda_{N}|}\end{pmatrix}\mathbf{R}^{-1}(\overline{\mathbf{q}}-\mathbf{q})$ (21b)
$=\mathbf{R}{\bm{\Lambda}_{\bf D}}(x;\Delta t)\mathbf{R}^{-1}(\overline{\mathbf{q}}-\mathbf{q})=\mathbf{D}(x;\Delta t)(\overline{\mathbf{q}}-\mathbf{q}),$ (21c)
@@
$d_{i}^{\text{RM}}(x;\Delta t)=\frac{\tilde{\Gamma}(x;\Delta t)}{|\lambda_{i}|}=\frac{1-\Gamma(x)}{\Delta t|\lambda_{i}|}.$ (22)
The inversely proportional relation between $d_{i}$ and $\Delta t$ is consistent with the increase in damping severity as the time step size decreases using the RM observed in the literature (see, e.g., [seng_slamming_2012]).
@@ 3.2 The RM with matrix-valued weight function (RM-M)
We propose a reframing of the RM such that it resembles the directional behavior of Karni’s damping far-field operator. We achieve this by allowing the weight function $\bm{\Gamma}$} to be matrix-valued and defined in terms of $\mathbf{R}$ (and therefore $\mathbf{q}$ in the nonlinear case). The process is described in Algorithm 2 and justified as follows. Proceeding analogously to the previous section, the relaxation step in Algorithm 2
$\mathbf{q}^{[n+1]}={{\bm{\Gamma}}}(x)\mathbf{q}^{*}+(\mathbf{I}-{\bm{\Gamma}}(x))\overline{\mathbf{q}},$ (23)
@@
$\partial_{t}\mathbf{q}=\tilde{{\bm{\Gamma}}}(x,\mathbf{q};\Delta t)(\overline{\mathbf{q}}-\mathbf{q}),$ (24)
@@
$\tilde{{\bm{\Gamma}}}(x,\mathbf{q};\Delta t)=\frac{1}{\Delta t}(\mathbf{I}-{{\bm{\Gamma}}}(x))=\mathbf{R}\begin{pmatrix}\frac{1-\Gamma_{1}(x)}{\Delta t}&0&\cdots&0\\ 0&\frac{1-\Gamma_{2}(x)}{\Delta t}&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&\frac{1-\Gamma_{N}(x)}{\Delta t}\end{pmatrix}\mathbf{R}^{-1}.$ (25)
@@
$d_{i}^{\text{RM-M}}(x)=\frac{1-\Gamma_{i}(x)}{\Delta t|\lambda_{i}|},$ (26)
@@
If $\mathbf{R}$ is independent of $\mathbf{q}$, the matrix-valued weight function $\bm{\Gamma}$} does not have to be computed at each time step in Algorithm 2. Additionally, if the far-field state $\overline{\mathbf{q}}$ is constant, we can define the translated function $\mathbf{v}=\mathbf{q}-\overline{\mathbf{q}}$ that satisfies the conservation law $\partial_{t}\mathbf{v}+\partial_{x}\mathbf{f}(\mathbf{v}+\overline{\mathbf{q}})=0$. This system may be solved with the far-field state $\bf 0$ and $\mathbf{q}$ recovered afterward. In this case, we have
[…]
@@
$\mathbf{q}^{*}\leftarrow$ Evolve a discretization of (1) over a time step $\Delta t$ with
 initial data $\mathbf{q}^{[n]}$ $\triangleright$ Hyperbolic step
@@
${\bm{\Lambda}}_{\bm{\Gamma}}(x)\leftarrowdiagonal$ matrix-valued function with entries $\Gamma_{i}(x)$
${\bm{\Gamma}}(x)\leftarrow\mathbf{R}{\bm{\Lambda}}_{\bm{\Gamma}}(x)\mathbf{R}^{-1}$ where $\mathbf{R}$ contains the right eigenvectors
 of $\mathbf{f}^{\prime}(\mathbf{q})$
$\mathbf{q}^{[n+1]}\leftarrow{\bm{\Gamma}}(x)\mathbf{q}^{*}+(\mathbf{I}-{\bm{\Gamma}}(x))\overline{\mathbf{q}}$ $\triangleright$ Relaxation step
@@
We illustrate the effect of damping waves corresponding to a particular characteristic field with a preliminary numerical experiment using the framework to be described in Section 5.1. An artificial setup where an incoming pulse is created inside the sponge layer is considered, as shown in Figure 2. The initial pulse is immediately damped by the scalar methods (RM and S-SDO), while just the right-going components of it are absorbed by the directional methods (RM-M and SDO). In a practical setting, the incoming pulse could correspond to, e.g., physical nonlinear interactions inside the sponge layer. While an artificially large pulse is used in Figure 2 for visualization purposes, such nonlinear interactions might not be less significant than spurious reflections due to the boundary treatment (see Section 5.3).
Figure 2: A Gaussian pulse is initialized inside the sponge layer to represent the effect of nonlinear interactions taking place there. Thus, we consider its left-going component as a physical wave to be preserved. The right-going sawtooth-like wave corresponds to disturbances introduced by the motion of a piston, which will be described in Section 4.1. The sponge layer is colored light grey. Above: Effect of directional (SDO) and scalar (S-SDO) damping operators applied to the incoming physical wave. Below: Effect of directional relaxation method with matrix-valued weight function (RM-M) and scalar relaxation method (RM) applied to the incoming physical wave.
@@ 3.3 On attaining higher order of accuracy in time with the RM
Let us consider Algorithm 2 for the linear problem (2), where the matrix-valued weight function $\bm{\Gamma}$ is independent of $\mathbf{q}$. Then, the relaxation step (23) could be written as a high-order discretization of a system $\partial_{t}\mathbf{q}=\mathcal{R}(g(\bm{\Gamma}(x),\Delta t))\,(\overline{\mathbf{q}}-\mathbf{q})$, where $\mathcal{R}$ is the stability function of some Runge-Kutta scheme [gottlieb_high_2009]. Moreover, we can interpret it as the exact solution to the problem
@@
${\bm{\Gamma}}(x;\Delta t)=\exp({-\Delta t\mathbf{W}(x;\Delta t)}).$ (29)
Since ${\bm{\Gamma}}=\mathbf{R}{\bm{\Lambda}}_{{\bm{\Gamma}}}\mathbf{R}^{-1}$, then $\mathbf{W}=\mathbf{R}{\bm{\Lambda}_{\bf W}}\mathbf{R}^{-1}$, where
@@
with $\Gamma_{i}$ and $w_{i}$ denoting the main diagonal entries of the matrices ${\bm{\Lambda}}_{\bm{\Gamma}}$ and $\bm{\Lambda}_{W}$} respectively. Thus, the $i$-th wave is damped at the rate
$d_{i}^{\mathbf{W}}(x)=\frac{w_{i}(x)}{|\lambda_{i}|}=\frac{-\ln\left(\Gamma_{i}(x)\right)}{\Delta t|\lambda_{i}|},$ (31)
and will be completely absorbed as $\Gamma_{i}\to 0$, i.e., at the boundary of the sponge layer $x_{\infty}$. This further justifies the approaches proposed by Seng [seng_slamming_2012] and Wasistho et al. [wasistho_simulation_1997] where, through different arguments, $\Gamma$ is substituted by $\Gamma^{C\Delta t}$ to reduce the damping sensitivity to the time step size.
@@
and a natural approach to get a higher-order accurate in time approximate solution to (32) would be to use the RM in the context of high-order operator splitting. For instance, we could use Strang splitting as described in Algorithm 3. However, in this setting, both splitting techniques are equivalent up to the choice of the scalar weight functions $\Gamma_{i}$. This can be seen as follows: Assuming that the solution is initialized as $\overline{\mathbf{q}}$ inside the sponge layer, the first relaxation step at $t^{[0]}=0$ does not alter the solution. Similarly, the second relaxation step in Algorithm 3 at the final time does not affect the solution in the computational domain, since it is local and no hyperbolic step is taken afterward. Finally, let $\mathbf{q}^{**}(t^{[n]})$ denote the second stage of Algorithm 3 with initial data $\mathbf{q}^{[n]}$, and let $\mathbf{q}^{*}(t^{[n+1]})$ denote the first stage of Algorithm 3 with initial data $\mathbf{q}^{[n+1]}$. Then, the following relation holds:
$\mathbf{q}^{*}(t^{[n+1]})$ $=\bm{\Gamma}\mathbf{q}^{[n+1]}+(\mathbf{I}-\bm{\Gamma})\overline{\mathbf{q}}$ (33a)
$=\bm{\Gamma}\left(\bm{\Gamma}\mathbf{q}^{**}(t^{[n]})+(\mathbf{I}-\bm{\Gamma})\overline{\mathbf{q}}\right)+(\mathbf{I}-\bm{\Gamma})\overline{\mathbf{q}}$ (33b)
$=\bm{\Gamma}^{2}\mathbf{q}^{**}(t^{[n]})+(\mathbf{I}-\bm{\Gamma}^{2})\overline{\mathbf{q}}.$ (33c)
@@
On the other hand, since $\bm{\Gamma}$ depends on $\mathbf{q}$ in the nonlinear case, the last observation is no longer valid, and Algorithms 2 and 3 are not equivalent a priori. The effect of applying the RM as in Algorithm 3 in the nonlinear setting is evaluated numerically in Section 5.3.
@@
Compute the weight function ${\bm{\Gamma}}(x)$ as in Algorithm 2
$\mathbf{q}^{*}\leftarrow{\bm{\Gamma}}(x)\mathbf{q}^{[n]}+(\mathbf{I}-{\bm{\Gamma}}(x))\overline{\mathbf{q}}$
@@
$\mathbf{q}^{**}\leftarrow$ Evolve a discretization of (1) over a time step $\Delta t$ with
 initial data $\mathbf{q}^{*}$ $\triangleright$ Hyperbolic step
$\mathbf{q}^{[n+1]}\leftarrow{\bm{\Gamma}}(x)\mathbf{q}^{**}+(\mathbf{I}-{\bm{\Gamma}}(x))\overline{\mathbf{q}}$
@@ 3.4 Nonlinear damping far-field operator (NDO)
We propose an alternative extension of the damping operator from Section 2.1 to hyperbolic systems of the form (1) with a general nonlinear flux function. Here, the change of variables $\mathbf{w}=\mathbf{R}^{-1}\mathbf{q}$ no longer allows us to decompose a nonlinear system of the form (1) into a system of scalar evolution equations. Instead, we consider the nonlinear characteristic variables $\bm{\alpha}=\mathbf{R}^{-1}\partial_{x}\mathbf{q}$ which, after some manipulations to System (1) (see [lefloch_hyperbolic_2002, Ch.1]), yield
@@
$\partial_{t}\alpha_{i}+\partial_{x}\left(\lambda_{i}(\mathbf{q})\alpha_{i}\right)=\mathcal{I}(\mathbf{q},\bm{\alpha}),\quad 1\leq i\leq N,$ (34)
@@
where $\mathcal{I}(\mathbf{q},\bm{\alpha})$ denotes the nonlinear interactions between the $N$ characteristic fields. Although this is not a decoupled system of equations, we can add a damping term at the right-hand side, resembling the modification done in (6) for the linear equations (2), to obtain
@@
$\partial_{t}\alpha_{i}+\partial_{x}\left(\lambda_{i}(\mathbf{q})\alpha_{i}\right)=\mathcal{I}(\mathbf{q},\bm{\alpha})-d_{i}(x)|\lambda_{i}(\mathbf{q})|\alpha_{i},\quad 1\leq i\leq N.$ (35)
@@
After multiplying this system from the left by $\mathbf{R}$ and integrating from $x$ to $\infty$ (inverting the process carried out to obtain (34)), we get
$\partial_{t}\mathbf{q}(t,x)+\partial_{x}\mathbf{f}(\mathbf{q}(t,x))$ $=\int_{x}^{+\infty}\mathbf{R}\bm{\Lambda}_{\mathbf{D}}(z){\mathbf{R}}^{-1}\partial_{z}\mathbf{q}(t,z)\,dz$ (36a)
$\approx\int_{x}^{x_{\infty}}\mathbf{D}(z,\mathbf{q})\partial_{z}\mathbf{q}(t,z)\,dz,$ (36b)
@@
where the artificial boundary at the end of the sponge layer is enforced through the assumption $\mathbf{q}(t,x)=\overline{\mathbf{q}}$ for $x\geq x_{\infty}$. If the damping rate functions $d_{i}$ are defined as in Section 2.1, the term (36b) vanishes inside of the computational domain $x<x_{s}$. Moreover, if the flux is linear and each $d_{i}$ is constant, the damping far-field operator of Section 2.1 is recovered.
@@
Some challenges that arise when dealing with the nonlinear damping term (36b) are its non-local nature and the presence of a nonconservative product. This term is significantly simplified if one can write $\mathbf{D}(z,\mathbf{q})=\partial_{\mathbf{q}}\mathbf{h}(z,\mathbf{q})$ for some function $\mathbf{h}\in C^{1}(\mathbb{R}\times\mathbb{R}^{N},\mathbb{R}^{N})$. Indeed, in that case, we would have
$\int_{x}^{x_{\infty}}\mathbf{D}(z,\mathbf{q})\partial_{z}\mathbf{q}(t,z)\,dz$ $=\int_{x}^{x_{\infty}}\partial_{\mathbf{q}}\mathbf{h}(z,\mathbf{q})\partial_{z}\mathbf{q}(t,z)\,dz$ (37a)
$=\int_{x}^{x_{\infty}}\frac{d}{dz}\mathbf{h}(z,\mathbf{q})-\partial_{z}\mathbf{h}(z,\mathbf{q})\,dz$ (37b)
$=\mathbf{h}(x_{\infty},\overline{\mathbf{q}})-\mathbf{h}(x,{\mathbf{q}(t,x)})-\int_{x}^{x_{\infty}}\partial_{z}\mathbf{h}(z,\mathbf{q}(t,z))\,dz.$ (37c)
@@
Nevertheless, the integrability conditions to guarantee the existence of such a function $\mathbf{h}$ do not hold for a general system and choice of $d_{i}$’s [papachristou_aspects_2019]. Thus, for the application studied in Section 5 we work with a discretization of the term (36b), described in Appendix C.